OpenStudy (anonymous):

The equation y=-1777x=27,153 can be used too predict the number y of gun deaths in the United Sates x years after 2000, that is x=0 corresponds to 2000, x=2 corresponds to 2002, x=4 corresponds to 2004, and so on. Preict the number of gun death in 200 and 2007. Inwhat year will the number of gun death be 11,160. 2004= 2007= what year will it be 11,160 Please help

OpenStudy (gw2011):

If I understand the problem correctly, here are the solutions: 2004: (-1777)(4)=-7108 27,153 - 7108 = 20,045 for the year 2004 2007: (-1777)(7)=-12,439 27,153 - 12,439 = 14,714 for the year 2007 The year for 11,160: y = -1777x = 27,153 27,153 - 1777x = 11,160 27,153 - 11,160 = 1777x 15,993 = 1777x 15,993/1777 = x 9 = x Therefore, this will be in the year 2009